Monitoring - A fix has been implemented and we are monitoring the results.
Jun 6, 16:19 BST
Identified - The issue has been identified and a fix is being implemented.
Jun 6, 16:16 BST
Investigating - We are currently investigating an issue where customers get an error when trying to pay by credit/debit card.
Jun 6, 16:15 BST

About This Site

Welcome to our public status page. Don't agree with what you see here? Email help@magicinternetgroup.com.

Servers Operational
90 days ago
100.0 % uptime
Today
Cloud804 Operational
90 days ago
100.0 % uptime
Today
AMS2 Operational
90 days ago
100.0 % uptime
Today
FRA1 Operational
90 days ago
100.0 % uptime
Today
LON1 Operational
90 days ago
100.0 % uptime
Today
NYC1 Operational
90 days ago
100.0 % uptime
Today
Customer Service Operational
Phone Support Operational
Live Chat Operational
Email/Ticket Operational
Help Centre Operational
Customer Portal Operational
Customer Website ? Operational
Billing Operational
Third party services Operational
SiteLock Operational
SEO Tools (marketgoo) Operational
Domains Operational
Registration Operational
Transfer Operational
DNS management Operational
Search Operational
Control Centre Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
had a major outage
had a partial outage
Past Incidents
Jun 6, 2020

Unresolved incident: Billing issues.

Jun 5, 2020

No incidents reported.

Jun 4, 2020

No incidents reported.

Jun 3, 2020

No incidents reported.

Jun 2, 2020

No incidents reported.

Jun 1, 2020

No incidents reported.

May 31, 2020

No incidents reported.

May 30, 2020

No incidents reported.

May 29, 2020

No incidents reported.

May 28, 2020

No incidents reported.

May 27, 2020

No incidents reported.

May 26, 2020

No incidents reported.

May 25, 2020

No incidents reported.

May 24, 2020

No incidents reported.

May 23, 2020
Resolved - Our engineering team has resolved the issue with networking in our NYC1 region. Networking should now be operating normally. If you continue to experience problems, please open a ticket with our support team. We apologize for any inconvenience.
May 23, 12:30 BST
Monitoring - Our engineering team has implemented a fix to resolve the networking issue in our NYC1 region and is monitoring the situation. We are monitoring the situation and will post an update as soon as the issue is fully resolved.
May 23, 12:17 BST
Identified - Our engineering team has identified the cause of the intermittent networking issue in our NYC1 region and is actively working on a fix. We will post an update as soon as additional information is available.
May 23, 11:29 BST